OpenStudy (anonymous):

somebody help plz! solving inequalities: -6/7y-6 is greater than or equal to 42 somebody help plz! solving inequalities: -6/7y-6 is greater than or equal to 42 @Mathematics

OpenStudy (anonymous):

is it -6/(7y-6)?

OpenStudy (anonymous):

-6/7y :its a fraction.

OpenStudy (anonymous):

yea I'm trying to figure out what is in the denominator of your fraction.

OpenStudy (anonymous):

\[-6/7y-6\ge42\]

OpenStudy (anonymous):

But is the y in the denominator? or is it 6/7 x y?

OpenStudy (anonymous):

-6/7 multiplied by y

jhonyy9 (jhonyy9):

(-6/7y) -6>= 42 -6-42y>=294y -6>=-252y y>= 6/252 y>= 1/42

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

wot?...y?????????????????